Circular No. 7121

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA IN THE SMALL MAGELLANIC CLOUD D. L. Welch, McMaster University, and the MACHO collaboration (cf.

IAUC 6312) report their discovery of an apparent nova at α = 1^{h}02^{m}21^{s}.8, δ = −73^{o}18′27′′ (equinox 2000.0), which peaked in brightness sometime between consecutive observations on 1998 Dec. 25.513 (V = 19.0) and 1999 Jan. 3.519 UT (V < 13.5; image saturated). Prior to eruption, a stellar object at the nova’s position had a mean magnitude of V = 19.0, with peak-to-peak brightness changes of 0.5 mag on a timescale of hundreds of days. On 1999 Mar. 3.463, it was at V ∼ 16, and it continues to fade.

XTE J1550–564 J. Homan, R. Wijnands, and M. van der Klis, Astronomical Institute

‘Anton Pannekoek’, University of Amsterdam, report: “The blackhole candidate x-ray transient source XTE J1550–564 (IAUC 7008) seems to have undergone a sudden major state transition. The source was observed with the RXTE PCA instrument on Mar. 4.80 UT. The source flux (2–20.0 keV) was 6.0 × 10^{−8} erg s^{−1} cm^{−2}. A quasiperiodic oscillation (QPO) was detected at 284.2 ± 1.3 Hz, with a width (FWHM) of 29 ± 4 Hz and an rms amplitude of 3.9 ± 0.2 percent (6.4–60 keV; 11.5σ). The QPO amplitudes in the energy ranges 2–60, 2–6.4, 6.4–13, and 13–60 keV, are 1.3 ± 0.2, < 0.8, 3.2 ± 0.2, and 8.4 ± 0.9 percent, respectively. This high-frequency QPO is most likely related to the 184-Hz QPO previously observed in XTE J1550–564 (IAUC 7025; Remillard et al. 1999, Ap.J. Let., submitted). Simultaneously with the high-frequency QPO, QPOs at 5.92 ± 0.02 Hz [FWHM 2.02 ± 0.05 Hz; rms amplitude 2.98 ± 0.03 percent (2-60 keV; 43.6σ)] and at 10.4 ± 0.1 Hz [FWHM 5.8 ± 0.3 Hz; rms amplitude 2.55 ± 0.06 percent (2–60 keV; 23.7σ)] are observed. An RXTE observation taken the day before (on Mar. 3.06), shows a 2–20.0-keV x-ray flux ∼ 20 percent lower than that of Mar. 4.80 (4.9 × 10^{−8} erg s^{−1} cm^{−2}), a softer x-ray spectrum, and a power spectrum with no QPO, dominated by a weak (2.7 ± 0.2 percent; 2-60 keV) power-law component with index −1.23 ± 0.04. These properties are typical for those of XTE J1550–564 during the last few months. This pronounced change in x-ray properties, which took place within 1.8 days, shows that XTE J1550–564 has undergone a major source state transition. Observations at all wavelengths are encouraged. More x-ray observations with the RXTE/PCA instrument are required to study the high-frequency QPO in more detail.”

1999 March 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green